# Boosting Adverse Weather Crowd Counting via Multi-queue Contrastive Learning

### v1

@@ Abstract
I Introduction
Crowd counting has attracted much attention recently due to its wide range of applications such as public safety, video surveillance, and traffic control. Currently, most of the crowd counting methods [1, 2, 3] are able to estimate the number of crowds well on the images recorded under normal weather conditions. However, when it comes to adverse weather conditions such as rain, haze, and snow, these methods face challenges in maintaining their performance due to domain differences and the limited availability of adverse weather images. For example, in the JHU-Crowd++ dataset [4], the number of images under rain conditions accounts for only 3% of the total dataset.
@@ I Introduction
To mitigate the negative influence of adverse weather, an intuitive remedy is to pre-process the images using image restoration modules before counting. Unfortunately, even when restoration modules can mitigate the obscuration caused by adverse weather, the restored images still exhibit significant domain differences from normal weather images. Moreover, the additionally introduced classification and enhancement modules significantly increase the computational burden. Training a separate model for each weather category is also not feasible because the limited samples of adverse weather conditions are insufficient to support the training. [5] introduces a transformer-based approach that addresses the issue of generating complementary information from image-specific degradation without the need for customized image enhancement modules. However, this transformer-based approach not only poses a heavy computational burden but also ignores the issue of weather class imbalance in the dataset.
The aim of this study is to enhance the robustness of crowd counting models in unknown adverse weather conditions while maintaining their performance under normal weather conditions without significantly increasing model complexity. To achieve this objective, we approach the problem of crowd counting, which includes both normal and adverse weather conditions, as an imbalanced multi-domain learning task. The key prerequisite for a model to tackle multi-domain learning is to have perception capabilities across different domains, meaning that the extracted information from each domain should possess discriminative characteristics. This aligns with the objective of contrastive learning [6, 7, 8]. Therefore, in this paper, we propose a two-stage method called Multi-queue Contrastive Learning (MQCL). This approach enables the backbone model to directly extract weather-aware representations, which are further refined by a refiner module. In the first stage, we use contrastive learning to distinguish the characteristics of different weather. However, since the class imbalance mentioned above, the gradient of the loss function of the vanilla contrastive method would be dominated by normal weather images, resulting in poor performance of representations and difficulty in refining and counting. To tackle such an imbalance problem, we design a simple yet effective contrastive learning method called multi-queue MoCo, which replaces the standard single queue in MoCo [7] with multiple queues, providing class-balanced key vectors. In the second stage, all normal weather samples in the queues are treated as positive samples of contrastive learning. Thus the refiner can be guided to convert the representations of adverse weather images to the domain of normal images. Benefiting from the effective representation learning, the refiner and the decoder can be designed to be light to meet our target of not significantly increasing the weight. Compared to the backbone model ConvNeXt [10, 11] used in this paper, our method only introduces 15.3% of the extra FLOPs and 12.7% parameters. Comparison of weight and accuracy of MQCL and state-of-the-art methods is shown in Fig. 1(a/b).
@@
II Related Work
II-A Crowd Counting under Normal Conditions
@@
Up to now, most single image crowd counting methods can be divided into two categories: density-map-based and localization-based crowd counting. Density-based methods aim to generate a density map, the pixel values of which represent the crowd density at that location. The sum of all pixel values in the density map is the estimated total number of the image. MCNN [1] is a pioneer in employing such a method. Benefiting from the multi-column design, MCNN can handle input images of arbitrary size or resolution. CSRNet [2] employs dilated CNN for the back end to deliver larger reception fields and to avoid pooling operations. More recently, Lin et al. proposed a multifaceted attention network [3] to improve transformer models in local spatial relation encoding. [21] redesigns the multi-scale neural network by introducing a hierarchical mixture of density experts. In addition to model architecture, loss function designing [13, 14] is also a focused area of regression-based crowd counting, which enables the models to effectively learn from ground truth.
@@
On the other hand, the target of detection-based methods is to predict head positions. Early methods obtain the positions by post processing from bounding boxes [31], density maps [37, 38]. However, these methods suffer some inevitable problems such as requirements for complex multi-scale representation and the preclusion of end-to-end training. More recently, Song et al. [15] proposed a model called P2PNet to directly predict a set of point proposals to represent heads in an image, being consistent with the human annotation results. They also proposed a new metric called density Normalized Average Precision (nAP) to provide more comprehensive and more precise performance evaluation. CLTR [33] is a concurrent work with P2PNet, introducing the Transformer architecture to the field of crowd localization and proposing a KMO-based Hungarian matcher to reduce the ambiguous points and generate more reasonable matching results.
II-B Crowd Counting under Adverse Conditions
@@
Existing deep-learning-based methods have achieved unprecedented success with crowd counting, but their performance degraded severely under adverse conditions (e.g., adverse weather) due to the disturbance to the brightness and gradient consistency. However, few research efforts have been made into this problem. Additional class conditioning blocks are utilized by [4] to augment the backbone module, which is trained via cross-entropy error using labels available in the dataset. [5] enables the model to extract weather information according to the degradation via learning adaptive query vectors, but the weight of the model is significantly increased due to the introduction of a Transformer-based module. [36] proposes a single-stage hazy-weather crowd counting method based on direction-aware attention aggregation. However, their method only focuses on the performance in hazy scenes and cannot handle various unknown weather conditions.
II-C Contrastive Learning
@@
Contrastive learning [8, 6] has attracted much attention due to its success in unsupervised representation learning. The target of it is to maximize the similarity of the representations between positive pairs while minimizing that of negative pairs. In recent years, there has been a lot of work to tap the potential of contrastive learning. For example, [7] builds a dynamic dictionary with a queue and a moving-averaged encoder to enable large-scale contrastive learning with dramatically low demand for memory. [9] investigates the contrastive loss and adapted contrastive learning to the field of supervised learning.
@@
Despite the significant success achieved by contrastive learning, Assran et al. [23] revealed that in the formulation of contrastive methods is an overlooked prior to learn features that enable uniform clustering of the data. This prior can hamper performance when pretraining on class-imbalanced data. To solve the class imbalance problem, Kang et al. proposed k-positive contrastive loss (KCL) [39], which draws k instances from the same class to form the positive sample set. Jiang et al. presented a framework called Model-Aware K-center (MAK) [40] to sample open-world unlabeled data to supplement samples for the tail-end classes. Balanced contrastive learning [41] tries to mitigate the unbalanced problem from the aspects of re-weighting the gradient and ensuring batch data balance. Targeted supervised contrastive learning (TSC) [43] improves the uniformity of the feature distribution by generating uniformly distributed cluster centers in advance. Hou et al. [42] clusters each head class into multiple subclasses of similar sizes as the tail classes and enforce representations to capture the two-layer class hierarchy between the original classes and their subclasses.
III Proposed Method
@@
Figure 2: The architecture of MQCL. The target of the WRL stage is to learn weather-aware representations via unsupervised contrastive learning. After the WRL stage, the supervised contrastive learning is utilized in the CRR stage to realize the refinement of the representations.
III-A Framework Overview
@@
The architecture of our method is illustrated in Fig. 2. As discussed earlier, we aim to directly enable the crowd counting backbone model to learn weather-aware feature representations and then refine it with a light-weight refiner. Thus, the decoder can focus on a single domain and generate high-quality density map, the sum of which is the estimated number. Since the refiner is premised on stable and consistent representation while it keeps evolving and is not stable during the representation learning stage, we divide the training into two stages to separate these two targets, namely Weather-aware Representation Learning (WRL) stage and Contrastive-learning-guided Representation Refining (CRR) stage, respectively.
@@
In the WRL stage, we use contrastive learning to enable the encoder to learn weather-aware representations. The contrastive learning strategy of the WRL stage is based on our proposed multi-queue MoCo, which can tackle the problem of weather class imbalance. The weights of the encoder and the decoder obtained during the WRL stage will be retained for the CRR stage. In the CRR stage, the encoder is freezed and the main task of this stage is to train a refiner. By treating the normal weather samples as positive examples of contrastive learning, the refiner is guided to pull the adverse weather representations towards normal weather domain. Finally, high-quality density map can be generated to realize precise counting.
III-B Weather-aware Representation Learning
@@
The model architecture of the WRL stage is illustrated in the upper half of Fig. 2. The contrastive learning method [8, 6] is utilized to train the encoders and endow it with the capability to extract weather-aware representations, i.e., images with similar weather conditions correspond to similar representations, whereas those with dissimilar conditions correspond to more distant representations.
@@
Our training structure is the same as MoCo [7], which consists of an encoder $\rm E_{Q}$, a momentum-updated encoder $\rm E_{K}$ and a decoder. The encoders $\rm E_{Q}$ and $\rm E_{K}$ extracts representations $\rm R_{Q}$ and $\rm R_{K}$, respectively, from different augmentations $\rm I_{Q}$ and $\rm I_{K}$ of an image. The representations $\rm R_{Q}$ and $\rm R_{K}$ will be further processed by our proposed multi-queue MoCo strategy, which tackles the problem of weather class imbalance and grant the representations with weather-aware nature, to obtain the contrastive loss. The details of multi-queue MoCo will be elaborated in section III-C. To ensure that the representation simultaneously contains crowd information, the target of the decoder is set as generating a density map under the supervision of the Bayesian loss [13] according to the representation $\rm R_{Q}$. The total loss of the WRL stage is:
@@
III-C Multi-queue MoCo
@@
III-D Contrastive-learning-guided Representation Refining
@@
The multiple queues obtained during the WRL stage have stored a sufficient number of vectors from various weather conditions. In this paper, we propose to preserve and freeze these multiple queues and employ contrastive learning to guide the refiner to convert the representation. In this stage, all the vectors in the normal-weather sub queue are treated as positive samples and those in other sub queues are treated as negatives. Since there is no longer a need to generate vectors K, the encoder $\rm E_{K}$ and projection head $\rm P_{K}$ are discarded. In order to maintain stable representations, the parameters of encoder $\rm E_{Q}$ are fixed at this stage. The projection head $\rm P_{Q}$ is preserved, fixed, and moved behind the refiner. We denote the vector projected from the refined representation as $\rm Q^{\prime}$. The model structure of the CRR stage is shown in the lower half of Fig. 2. The loss function of contrastive learning in this stage is calculated as follows:
$\mathcal{L}_{contra2}=\sum_{i\in I}\frac{-1}{|\mathbf{N}|}\sum_{p\in\mathbf{N}}\log\frac{\exp(Q^{\prime}_{i}\cdot{\rm K_{p}}/\tau)}{\sum_{a\in\mathbf{A}}\exp(R({\rm Q_{i}})\cdot{\rm K_{a}}/\tau)},$ (3)
@@
IV Experiments and Discussions
IV-A Experiment Setup
The ConvNeXt [10] is employed as our backbone model and we use the ConvNeXt-T version for simplicity. The structure of ConvNeXt-T is: C = (96, 192, 384, 768), B = (3, 3, 9, 3), where C stands for the number of channels and B is the number of ConvNeXt blocks in each stage. Following [11], we replace the linear layer at the end of the ConvNeXt by an upsampling block to keep the downsampling rate at 8. The encoder $\rm E_{Q}$ and encoder $\rm E_{K}$ both have the same structure as the first three stages of ConvNeXt and the decoder corresponds to the final stage and the upsampling module. The pre-trained weights of ImageNet-22K [24] are loaded as the initial parameters. Two-layer MLPs are employed in the projection heads, the output dimension of which is 2048 and 128. The refiner consists of three ConvNeXt blocks with input and output dimensions of 768, initialized with random parameters.
@@ IV-A Experiment Setup
Following the convention of existing works [2, 3], we adopt Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the metrics to evaluate the methods.
TABLE I: Quantitative results comparing with the state-of-the-art methods on the NWPU-Weather dataset. The numbers in parentheses represent the rankings of our method.
@@
IV-B NWPU-Weather Dataset
Consistent with the JHU-Crowd++ dataset, our NWPU-Weather dataset keeps the imbalance of weather types, the hazy and rainy scenes are synthesized by a probability of 5%, respectively. The distribution of the weather conditions in the dataset is shown in Table II. We follow the approach outlined in [26] to synthesize the weather scenarios. The intensity, density, and angle of the rain are set as random values. The depth maps required during the haze synthesis process are estimated by ZoeDepth [27] and the intensity of the haze is also random. Examples of the images in the NWPU-Weather images are shown in the Fig. 4.
TABLE II: The distribution of the number of images under different weather conditions in the NWPU-Weather dataset.
Stage Normal Haze Rain
Train 2365 120 124
Val 446 31 23
Test 442 28 30
Total 3253 179 177
Except for the MQCL, we also benchmark several representative counting networks, providing an overview of the state-of-the-art performance, including CSRNet [2], DM-Count [28], KDMG [29], ConvNeXt [10] and MAN [3]. Note that since the model structure of AWCC-Net [5] is not fully elucidated and the code is not available at present, we do not provide the performance. Following the convention of existing works [2, 3], we adopt Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the metrics to evaluate the methods. The quantitative results of counting accuracy are listed in Table I. Compared with the baseline ConvNeXt, MQCL exhibits a significant performance improvement under adverse weather conditions. MAE and RMSE are decreased by 13.8% and 12.6%, suggesting that the proposed representation learning and refining strategy do indeed boost the robustness of the model under adverse weather conditions. Moreover, MQCL can also improve the performance under normal conditions, decreasing MAE and RMSE by 11.1% and 7.0%, respectively. We believe that this can be attributed to the contrastive learning strategy employed in this paper, which considers augmentations from the same image as positive examples. This strategy not only assists the model in weather perception but also strengthens the model’s ability to recognize different scenes. Sampling from the same image ensures that the positive examples not only share the same weather conditions but also possess similar scene characteristics. Compared with the previously best method MAN, MQCL has also achieved significant performance improvement. The MAE/RMSE under normal and adverse weather conditions have reduced by 3.9%/5.2% and 11.9%/5.3%, respectively.
TABLE III: Quantitative results comparing with the state-of-the-art methods on the JHU-Crowd++ dataset. The numbers in parentheses represent the rankings of our method.
@@ IV-B NWPU-Weather Dataset
TABLE IV: Comparison of computational complexity and the number of parameters. The computational complexity is measured by FLOPs when inferencing images with the size of 384 $\times$ 384.

### v2

@@ Abstract
journal: Pattern Recognition
[…]
[a1]organization=School of Computer Science and Engineering, Nanjing University of Science and Technology,city=Nanjing, postcode=210094, state=Jiangsu, country=China
[…]
[a2]organization=Guangxi Key Lab of Multi-source Information Mining & Security, Guangxi Normal University,city=Guilin, postcode=541004, state=Guangxi, country=China
1 Introduction
Crowd counting has attracted much attention recently due to its wide range of applications such as public safety, video surveillance, and traffic control. Currently, most of the crowd counting methods [45; 19; 22] are able to estimate the number of crowds well on the images recorded under normal weather conditions. However, when it comes to adverse weather conditions such as rain, haze, and snow, these methods face challenges in maintaining their performance due to domain differences and the limited availability of adverse weather images. What’s worse, labeled adverse weather crowd images are exceedingly rare. For example, in the JHU-Crowd++ dataset [34], the number of images under rain conditions accounts for only 3% of the total dataset.
To mitigate the negative influence of adverse weather, an intuitive remedy is to pre-process the images using image restoration modules before counting. Unfortunately, even when restoration modules can mitigate the obscuration caused by adverse weather, the restored images still exhibit significant domain differences from normal weather images. Moreover, the additionally introduced classification and enhancement modules significantly increase the computational burden. Training a separate model for each weather category is also not feasible because the limited samples of adverse weather conditions are insufficient to support the training. Huang et al. [10] introduces a transformer-based approach that addresses the issue of generating complementary information from image-specific degradation without the need for customized image enhancement modules. However, this transformer-based approach not only poses a heavy computational burden but also ignores the issue of weather class imbalance in the dataset.
The aim of this study is to enhance the robustness of crowd counting models in unknown adverse weather conditions while maintaining their performance under normal weather conditions without significantly increasing model complexity. To achieve this objective, we approach the problem of crowd counting, which includes both normal and adverse weather conditions, as an imbalanced multi-domain learning task. The key prerequisite for a model to tackle multi-domain learning is to have perception capabilities across different domains, meaning that the extracted information from each domain should possess discriminative characteristics. This aligns with the objective of contrastive learning [3; 8; 30]. Therefore, in this paper, we propose a two-stage method called Multi-queue Contrastive Learning (MQCL). This approach enables the backbone model to directly extract weather-aware representations, which are further refined by a refiner module. In the first stage, we use contrastive learning to distinguish the characteristics of different weather. However, since the class imbalance mentioned above, the gradient of the loss function of the vanilla contrastive method would be dominated by normal weather images, resulting in poor performance of representations and difficulty in refining and counting. To tackle such an imbalance problem, we design a simple yet effective contrastive learning method called multi-queue MoCo, which replaces the standard single queue in MoCo [8] with multiple queues, providing class-balanced key vectors. In the second stage, all normal weather samples in the queues are treated as positive samples of contrastive learning. Thus the refiner can be guided to convert the representations of adverse weather images to the domain of normal images. Benefiting from the effective representation learning, the refiner and the decoder can be designed to be light to meet our target of not significantly increasing the weight. Comparison of weight and accuracy of MQCL and state-of-the-art methods is shown in Fig. 1(a/b).
@@ 1 Introduction
2 Related Work
2.1 Crowd Counting under Normal Conditions
@@
Up to now, most single image crowd counting methods can be divided into two categories: density-map-based and localization-based crowd counting. Density-based methods aim to generate a density map, the pixel values of which represent the crowd density at that location. The sum of all pixel values in the density map is the estimated total number of the image. MCNN [45] is a pioneer in employing such a method. Benefiting from the multi-column design, MCNN can handle input images of arbitrary size or resolution. CSRNet [19] employs dilated CNN for the back end to deliver larger reception fields and to avoid pooling operations. More recently, Lin et al. proposed a multifaceted attention network [22] to improve transformer models in local spatial relation encoding. Du et al. [6] redesigns the multi-scale neural network by introducing a hierarchical mixture of density experts.
@@
On the other hand, the target of localization-based methods is to predict head positions. Song et al. [35] propose a model called P2PNet to directly predict a set of point proposals to represent heads in an image, being consistent with the human annotation results. They also proposed a new metric called density Normalized Average Precision (nAP) to provide more comprehensive and more precise performance evaluation. CLTR [21] is a concurrent work with P2PNet, introducing the Transformer architecture to the field of crowd localization and proposing a KMO-based Hungarian matcher to reduce the ambiguous points and generate more reasonable matching results.
2.2 Crowd Counting under Adverse Conditions
@@
Existing deep-learning-based methods have achieved unprecedented success with crowd counting, but their performance degraded severely under adverse conditions (e.g., adverse weather) due to the disturbance to the brightness and gradient consistency. However, few research efforts have been made into this problem. Additional class conditioning blocks are utilized by [34] to augment the backbone module, which is trained via cross-entropy error using labels available in the dataset. AWCC [10] enables the model to extract weather information according to the degradation via learning adaptive query vectors, but the weight of the model is significantly increased due to the introduction of a Transformer-based module. Kong et al. [15] proposes a single-stage hazy-weather crowd counting method based on direction-aware attention aggregation. However, their method only focuses on the performance in hazy scenes and cannot handle various unknown weather conditions.
2.3 Contrastive Learning
@@
Contrastive learning [30; 3] has attracted much attention due to its success in unsupervised representation learning. The target of it is to maximize the similarity of the representations between positive pairs while minimizing that of negative pairs. He et al. [8] builds a dynamic dictionary with a queue and a moving-averaged encoder to enable large-scale contrastive learning with dramatically low demand for memory. [14] investigates the contrastive loss and adapted contrastive learning to the field of supervised learning.
@@
Despite the significant success achieved by contrastive learning, Assran et al. [1] revealed that in the formulation of contrastive methods is an overlooked prior to learn features that enable uniform clustering of the data. This prior can hamper performance when pretraining on class-imbalanced data. To solve the class imbalance problem, Kang et al. proposed k-positive contrastive loss (KCL) [13], which draws k instances from the same class to form the positive sample set. Balanced contrastive learning [46] tries to mitigate the unbalanced problem from the aspects of re-weighting the gradient and ensuring batch data balance. Hou et al. [9] clusters each head class into multiple subclasses of similar sizes as the tail classes and enforce representations to capture the two-layer class hierarchy between the original classes and their subclasses.
3 Proposed Method
@@
Figure 2: The architecture of MQCL. The target of the WRL stage is to learn weather-aware representations via unsupervised contrastive learning. After the WRL stage, a refiner is trained in the CRR stage to pull the adverse weather representations towards normal weather domain.
3.1 Framework Overview
@@
The architecture of our method is illustrated in Fig. 2. As discussed earlier, we aim to directly enable the crowd counting backbone model to learn weather-aware feature representations and then refine it with a light-weight refiner. Thus, the decoder can focus on a single domain and generate high-quality density map, the sum of which is the estimated number. Since the refiner is premised on stable and consistent representation while it keeps evolving and is not stable during the representation learning stage, we divide the training into two stages to separate these two targets, namely Weather-aware Representation Learning (WRL) stage and Contrastive-learning-guided Representation Refining (CRR) stage, respectively. After training is completed, the model of the CRR stage is used for inference.
@@
In the WRL stage, contrastive learning is utilized to enable the encoder to learn weather-aware representations. The contrastive learning strategy of the WRL stage is based on our proposed multi-queue MoCo, which can tackle the problem of weather class imbalance. The weights of the encoder, the decoder and the queues obtained during the WRL stage will be retained for the CRR stage. In the CRR stage, the encoder is freezed and the main task of this stage is to train a refiner. By treating the normal weather samples as positive examples of contrastive learning, the refiner is guided to pull the adverse weather representations towards normal weather domain. Finally, high-quality density map can be generated to realize precise counting. Note that the tiny version of ConvNeXt [26] is used as the encoder in our work. In terms of the decoder’s structure, we remain consistent with MFA [23], which primarily consists of two transposed convolutional layers.
3.2 Weather-aware Representation Learning
@@
The model architecture of the WRL stage is illustrated in the upper half of Fig. 2. The contrastive learning method is utilized to train the encoders and endow it with the capability to extract weather-aware representations, i.e., images with similar weather conditions correspond to similar representations, whereas those with dissimilar conditions correspond to more distant representations.
@@
Our training structure is the same as MoCo [8], which consists of an encoder $\rm E_{Q}$, a momentum-updated encoder $\rm E_{K}$ and a decoder. The encoders $\rm E_{Q}$ and $\rm E_{K}$ extracts representations $\rm R_{Q}$ and $\rm R_{K}$, respectively, from different augmentations $\rm I_{Q}$ and $\rm I_{K}$ of an image. The representations $\rm R_{Q}$ and $\rm R_{K}$ will be further processed by our proposed multi-queue MoCo strategy, which tackles the problem of weather class imbalance and grant the representations with weather-aware nature, to obtain the contrastive loss. The details of multi-queue MoCo will be elaborated in Section 3.3. To ensure that the representation simultaneously contains crowd information, the target of the decoder is set as generating a density map under the supervision of the Bayesian loss [28] according to the representation $\rm R_{Q}$. The total loss of the WRL stage is:
@@
3.3 Multi-queue MoCo
@@
3.4 Contrastive-learning-guided Representation Refining
@@
The multiple queues obtained during the WRL stage have stored a sufficient number of vectors from various weather conditions. In this paper, we propose to preserve and freeze these queues and employ contrastive learning to guide the refiner to convert the representation. In this stage, all the vectors in the normal-weather sub queue are treated as positive samples and those in other sub queues are treated as negatives. Since there is no longer a need to generate vectors K, the encoder $\rm E_{K}$ and projection head $\rm P_{K}$ are discarded. In order to maintain stable representations, the parameters of encoder $\rm E_{Q}$ are fixed at this stage. The projection head $\rm P_{Q}$ is preserved, fixed, and moved behind the refiner. We denote the vector projected from the refined representation as $\rm Q^{\prime}$. The model structure of the CRR stage is shown in the lower half of Fig. 2. The loss function of contrastive learning in this stage is calculated as follows:
$\mathcal{L}_{contra2}=\sum_{i\in I}\frac{-1}{|\mathbf{N}|}\sum_{p\in\mathbf{N}}\log\frac{\exp(\rm Q^{\prime}_{i}\cdot{\rm K_{p}}/\tau)}{\sum_{a\in\mathbf{A}}\exp({\rm Q^{\prime}_{i}}\cdot{\rm K_{a}}/\tau)},$ (3)
@@
4 Experiments and Discussions
4.1 Experiment Setup
The ConvNeXt [26] is employed as the encoder and we use the tiny version for simplicity. The structure of ConvNeXt-T is: C = (96, 192, 384, 768), B = (3, 3, 9, 3), where C stands for the number of channels and B is the number of ConvNeXt blocks in each stage. The decoder’s structure is consistent with MFA [23], which is an upsampling module composed of two transposed convolutional layers. The encoder $\rm E_{Q}$ and encoder $\rm E_{K}$ both have the same structure as ConvNeXt-T (the classification head is discarded) and the structure of the decoder keeps consistent with MFA [23], which primarily consists of two transposed convolutional layers. The pre-trained weights of ImageNet-22K [31] are loaded as the initial parameters. Two-layer MLPs are employed in the projection heads, the output dimension of which is 2048 and 128. The refiner consists of three ConvNeXt blocks with input and output dimensions of 768, initialized with random parameters.
@@ 4.1 Experiment Setup
Following the convention of existing works [19; 22], we adopt Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the metrics to evaluate the methods.
Table 1: The distribution of the number of images under different weather conditions in the NWPU-Weather dataset.
Stage Normal Haze Rain
Train 2365 120 124
Val 446 31 23
Test 442 28 30
Total 3253 179 177
4.2 NWPU-Weather Dataset
Table 2: Quantitative results comparing with the state-of-the-art methods on the NWPU-Weather dataset. The numbers in parentheses represent the rankings of our method.
@@ 4.2 NWPU-Weather Dataset
Consistent with the JHU-Crowd++ dataset, our NWPU-Weather dataset keeps the imbalance of weather types, the hazy and rainy scenes are synthesized by a probability of 5%, respectively. The distribution of the weather conditions in the dataset is shown in Table 1. We follow the approach outlined in [17] to synthesize the weather scenarios. The intensity, density, and angle of the rain are set as random values. The depth maps required during the haze synthesis process are estimated by ZoeDepth [2] and the intensity of the haze is also random. Examples of the images in the NWPU-Weather images are shown in the Fig. 4.
Except for the MQCL, we also benchmark several representative counting networks, providing an overview of the state-of-the-art performance. Note that since the model structure of AWCC-Net [10] is not fully elucidated and the code for training is not available, we do not provide the performance. The structure our baseline ConvNeXt in Table 3 is consistent of the original model in the work [26], and we only replace the classification head with the same decoder used in MQCL. The quantitative results of counting accuracy are listed in Table 2. Compared with the baseline ConvNeXt, MQCL exhibits a significant performance improvement under adverse weather conditions. MAE and RMSE are decreased by 13.8% and 12.6%, suggesting that the proposed representation learning and refining strategy do indeed boost the robustness of the model under adverse weather conditions. Moreover, MQCL can also improve the performance under normal conditions, decreasing MAE and RMSE by 11.1% and 7.0%, respectively. We believe that this can be attributed to the contrastive learning strategy employed in this paper, which considers augmentations from the same image as positive examples. This strategy not only assists the model in weather perception but also strengthens the model’s ability to recognize different scenes. Sampling from the same image ensures that the positive examples not only share the same weather conditions but also possess similar scene characteristics. Compared with the previously best method MAN, MQCL has also achieved significant performance improvement. The MAE/RMSE under normal and adverse weather conditions have reduced by 3.9%/5.2% and 11.9%/5.3%, respectively.
Table 3: Quantitative results comparing with the state-of-the-art methods on the JHU-Crowd++ dataset. The numbers in parentheses represent the rankings of our method.
@@
4.3 JHU-Crowd++ Dataset
Table 4: Comparison of computational complexity and the number of parameters. The computational complexity is measured by FLOPs when inferencing images with the size of 384 $\times$ 384.